OpenStudy (anonymous):

verify the identity... tan(theta)=cos(theta)/sin(theta) x cot^2

OpenStudy (zzr0ck3r):

\(\frac{\cos(\theta)}{\sin(\theta)\cot^2(\theta)}=\frac{\cos(\theta)}{\sin(\theta)(\frac{\cos(\theta)}{\sin(\theta)})^2}=\frac{\cos(\theta)}{\sin(\theta)\frac{\cos^2(\theta)}{\sin^2(\theta)})}=\frac{\cos(\theta)}{\cancel{\sin(\theta)}\frac{\cos^2(\theta)}{\sin^{\cancel{2}}(\theta)}}=\frac{\cos(\theta)}{\frac{\cos^2(\theta)}{\sin(\theta)}}=\cancel{\cos(\theta)}\frac{\sin(\theta)}{\cos^{\cancel{2}}(\theta)}\\=\frac{\sin(\theta)}{\cos(\theta)}=\tan(\theta) \ \ \ \ \ \ \ \ \ \ \ \ \ _\square\)

OpenStudy (zzr0ck3r):

I know, beautiful right?

OpenStudy (zzr0ck3r):

jk:0